ThreePoint Sports Inc. manufactures basketballs for the Women’s National Basketball Association (WNBA). For the first 6 months of 2020, the company reported the following operating results while operating at 80% of plant capacity and producing 118,800 units.

Amount
Sales $4,514,400
Cost of goods sold 3,633,000
Selling and administrative expenses 517,540
Net income $363,860


Fixed costs for the period were cost of goods sold $960,000, and selling and administrative expenses $274,000.

In July, normally a slack manufacturing month, ThreePoint Sports receives a special order for 10,000 basketballs at $29 each from the Greek Basketball Association (GBA). Acceptance of the order would increase variable selling and administrative expenses $0.77 per unit because of shipping costs but would not increase fixed costs and expenses.

Part 1

(a) Prepare an incremental analysis for the special order. (Round all per unit computations to 2 decimal places, e.g. 15.25. Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Reject
Order
Accept
Order
Net Income
Increase
(Decrease)
Revenues $enter revenues in dollars $enter revenues in dollars $enter revenues in dollars
Cost of goods sold enter the cost of goods sold in dollars enter the cost of goods sold in dollars enter the cost of goods sold in dollars
Selling and administrative expenses enter selling and administrative expenses in dollars enter selling and administrative expenses in dollars enter selling and administrative expenses in dollars
Net income $enter net income in dollars $enter net income in dollars $enter net income in dollars



(b) Should ThreePoint Sports Inc. accept the special order?

select between Yes and NoYesNo YesNo

eTextbook and Media

  • eTextbook

Part 2

What is the minimum selling price on the special order to produce net income of $5.11 per ball? (Round answer to 2 decimal places, e.g. 15.25.)

Minimum selling price $enter the minimum selling price rounded to 2 decimal places

eTextbook and Media

  • eTextbook
Answer #1

Variable cost of goods sold per unit = (3633000-960000)/118800 = $22.5
Variable selling and administrative expenses =(517540-274000)/118800= $2.05
For special order = $2.05+$0.77 = $2.82

Reject Order Accept Order Net Income Increase (Decrease)
Revenues $                  -   $      2,90,000 $       2,90,000 =10000*29
Cost of Goods Sold $                  -   $      2,25,000 $      -2,25,000 =10000*22.5
Selling and Administrative expenses $                  -   $          28,200 $         -28,200 =10000*2.82
Net Income $                  -   $          36,800 $          36,800

(b)
company should accept the special order, as it results in Increase in net income by $36800

2.

Cost of Goods Sold $      2,25,000
Selling and Administrative expenses $          28,200
Total Expenses $      2,53,200
Units 10000
Cost per unit $            25.32
Net Income $              5.11
Minimum Selling Price $            30.43